OpenStudy (anonymous):

Write an equation of the line containing the given point and perpendicular to the given line (6, -5); 4x+9y=7 I have tried this problem 2 times it took forever and I got it wrong both times.

OpenStudy (anonymous):

If anyone can help me on that one, i also need to know (-2,3); 7x=5y+3

OpenStudy (anonymous):

I lied I have another one too: (3,7); x+4y=9

OpenStudy (anonymous):

I need to know these in the next hour.

OpenStudy (anonymous):

For the first question, two lines are perpendicular when their slopes are negative reciprocals (that is to say slope1 times slope2 = -1). So if you solve the given (4x+9y=7) line in terms of y, you get: y = (4/9)x-(7/9). This means the slope of that line is 4/9. So, the slope of the line you are trying to find would be -9/4. Now use this slope and the given point [(x0,y0)=(6,-5)] in the equation y-y0=m(x-x0) where m is your slope.

OpenStudy (anonymous):

Okay that is what I did but i did not use a negative i kept -9/4 positive

OpenStudy (anonymous):

Same process for the other two questions. Solve for y in the given equation, the coefficient of your x is your slope, sub that and your given point into the equation y-y0=m(x-x0) and solve for y again.

OpenStudy (anonymous):

I am just horrible at math, I swear what you are saying sounds so confusing to me

OpenStudy (anonymous):

Sorry, my mistake. The slope of the given line would be -4/9 so the slope you want to use is 9/4 as you said.

OpenStudy (anonymous):

o okay wow. i am so lost, I have this assignment due tonight, and I need like 6 more questions to do, ive worked on them all morning and most of the night and had a family memeber help, I still have no idea what im doing